OpenStudy (anonymous):

Simplify and leave your answer as a proper fraction or a mixed number. Leave a space between the whole number and the fraction part: 15 and 3/10 - 6 and 3/8. Enter your answer as a reduced mixed number or as a whole number.

OpenStudy (ajprincess):

\[15 \frac{ 3 }{ 10 }-6\frac{ 3 }{ 8 }\] First let us make these mixed fractions as improper fractions. \[15\frac{ 3 }{ 10 }\] Multiply 15 by 10. 15*10=150 Add it with 3. 150+3=153 Write this as numerator and 10 as the denominator. \[\frac{ 153 }{ 10}\] this is the equivalent improper fraction to the mixed number \[15\frac{ 3 }{ 10 }\]. Similarly can u convert [6\frac{ 3 }{ 8 }\] as improper fraction? @Rikiamyers

OpenStudy (anonymous):

i don't know

OpenStudy (ajprincess):

Well do u understand upto vat I have typed?

OpenStudy (anonymous):

Yes but I don't know how to do Similarly can u convert [6\frac{ 3 }{ 8 }\] as improper fraction?

OpenStudy (ajprincess):

\[6\frac{ 3 }{ 8 }\] can u convert this as improper fraction?

OpenStudy (anonymous):

no

OpenStudy (ajprincess):

 This how it is done. Do u get it?